OpenStudy (12wishs):

Please help Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Number of years 1 2 3 Option 1 (amount in dollars) 1100 -1210-1331 Option 2 (amount in dollars) 1100 -1200-1300

OpenStudy (welshfella):

Hint: Option 1 divide the yesr 2 value by the year 1 value and the year 3 by the year 2 value and you'll find a common ratio.

OpenStudy (12wishs):

0.91?

OpenStudy (welshfella):

No year 2 / year 1 year 3 / yaer 2

OpenStudy (12wishs):

0.90

OpenStudy (welshfella):

1210/1100 = 0.90?

OpenStudy (12wishs):

no it equals 1.10

OpenStudy (welshfella):

thats more like it

OpenStudy (welshfella):

so starting with 1100 you multiply each number by 1.1 to get the next number

OpenStudy (12wishs):

ya I tried that but in the function I made f(x)=1100(1.10)^x it doesn't work.

OpenStudy (welshfella):

that is incorrect because for example if you put x ( or n is best) = 1 then the YEAR 1 AMOUNT IS 1210 not 1100 so instead of n being the exponent what do you think you need? For year n = 1 you want the function to have the value 1100(1.1)^0 which kives 1100.

OpenStudy (welshfella):

* gives

OpenStudy (welshfella):

Th exponent must be 'n - what'?

OpenStudy (12wishs):

I honestly don't know

OpenStudy (welshfella):

For n = 1, n - what has to be 0?

OpenStudy (welshfella):

let the 'what' be x n - x = 0 if n = 1 1 - x = 0 x = ?

OpenStudy (12wishs):

1

OpenStudy (welshfella):

exactly so the exponent is (n -1) and the formula is f(n) = 1100(1.1)^(n-1)

OpenStudy (12wishs):

wow it works thank you!

OpenStudy (welshfella):

yw

OpenStudy (12wishs):

can you help with option 2 as well?

OpenStudy (welshfella):

For option 2 there is a COMMON DIFFERENCE of 100 and the function wiil require 1100 to be part of it and you need to ADD 100 times an expression in n.

OpenStudy (welshfella):

No exponent here - you multuply.

OpenStudy (welshfella):

f(n) = 1100 + 100(something)

OpenStudy (12wishs):

1100*1.091=1200.1 this is as close as I got

OpenStudy (welshfella):

no - just think what the 'something' might be

OpenStudy (12wishs):

ok I've thought about it and I've come to the conclusion that its zero.

OpenStudy (welshfella):

Its zero for year 1

OpenStudy (welshfella):

This is similar to Otion 1

OpenStudy (welshfella):

*Option 1.

OpenStudy (12wishs):

is it one?

OpenStudy (welshfella):

No- you forget that an 'n' is requred in the function?

OpenStudy (welshfella):

an expression containig n.

OpenStudy (12wishs):

oh wait n would be 1 for the second year ,2 for the third year and so on and so forth

OpenStudy (welshfella):

yes

OpenStudy (12wishs):

ok I get it now thanks

OpenStudy (welshfella):

- not n - but an expression containing n

OpenStudy (12wishs):

n-1?

OpenStudy (welshfella):

the 'something in the parentheses is n-1 again.

OpenStudy (welshfella):

yep

OpenStudy (12wishs):

ok so like for year one n would be 1-1 and then for year 2 it would be 2-1 right?

OpenStudy (welshfella):

gtg hope you have learnt something - i think you have..

OpenStudy (welshfella):

exactly

OpenStudy (welshfella):

gtg right now

OpenStudy (12wishs):

thanks again have a good rest of the day